OpenStudy (anonymous):

A value grows exponentially each year, the value after four years and ten years has grown to 2385 and 3106. The percentage value grows each year? ??

OpenStudy (kropot72):

2385 = A(1 + r)^4 3106 = A(1 + r)^10 Where A is the original amount, r is the annual interest rate and the exponents are the number of years Taking logs of both sides of the above equations gives: \[\ln 2385=\ln A + 4\ln (1+r).............(1)\] \[\ln 3106=\ln A+10\ln (1+r).............(2)\] Subtracting (1) from (2) gives: \[\ln 3106-\ln 2385=6\ln (1+r)\] From which we get\[\ln (1+r)=0.44\] Can you now solve for r?

OpenStudy (kropot72):

\[\ln (1+r)=0.44\] Therefore\[1+r=e ^{0.44}=1.045\] \[r=1.045-1=?\]

OpenStudy (anonymous):

r = 0,045? is that the % ? , why 1+r=e0.44 ? dont understand where the e comes from?

OpenStudy (kropot72):

The exponential number e is the base of natural logarithms. The percentage rate is 0.045 * 100 = 4.5%

OpenStudy (anonymous):

Thanks :) I always have to use e when i use Ln ?

OpenStudy (kropot72):

You're welcome :) From the basic laws of logarithms, if\[e ^{a}=b\] then\[\ln b=a\]